Jefferson Lab HALL A Logbook for December 00 - All EXCEPT Automatic Entries

  • Main INDEX

    Friday

    
    

    Tuesday

    
    

    Wednesday

  • 53888 12/20/00 12:14 Eugene Update for Moller from Dec,14
  • Tuesday

  • 53887 12/19/00 10:25 Jack Gas mixing system
  • 53885 12/19/00 07:52 milbrath target lifter problems
  • 53884 12/19/00 07:07 R. Michaels access to swap IOCs, check target
  • 53883 12/19/00 06:45 Saha Scanner 1 Run
  • 53882 12/19/00 06:41 R. Michaels a run at intermediate rate
  • 53878 12/19/00 06:09 R. Michaels deadtime study
  • 53869 12/19/00 04:50 R. Michaels FPP trip
  • 53862 12/19/00 02:55 R. Michaels correction: position still -1, zero it and re-run
  • 53859 12/19/00 02:49 R. Michaels 15 cm dummy target run
  • 53858 12/19/00 02:44 R. Michaels beam scan with optics target
  • 53853 12/19/00 02:18 R. Michaels X = -1mm, HARP scan
  • 53850 12/19/00 02:04 R. Michaels HARP scan at x = 3.8
  • 53839 12/19/00 01:11 R. Michaels HARP run 1171
  • 53835 12/19/00 00:44 R. Michaels special runs
  • 53829 12/19/00 00:00 LeRose Shift Summary
  • Monday

  • 53828 12/18/00 23:36 LeRose FPP Trip
  • 53825 12/18/00 22:55 LeRose Front FPP trip
  • 53820 12/18/00 21:50 LeRose Half wave plate in
  • 53809 12/18/00 18:41 Eugene Moller results
  • 53804 12/18/00 16:03 Ole Hansen Shift Summary
  • 53792 12/18/00 08:02 R. Michaels owl shift summary
  • 53789 12/18/00 06:49 R. Michaels half waveplate change (after run1978)
  • 53788 12/18/00 06:43 R. Michaels magnet screen (2nd try)
  • 53784 12/18/00 06:35 R. Michaels beam stability (typical plot)
  • 53783 12/18/00 06:35 R. Michaels magnet screen
  • 53776 12/18/00 04:00 R. Michaels Calo oddity -- ADC 9 showed up
  • 53773 12/18/00 02:43 R. Michaels No comparison (kumac problem ?)
  • 53770 12/18/00 01:47 Neyret Compton Beam polarization
  • Sunday

  • 53765 12/17/00 23:51 S. Dieterich Shift summary
  • 53745 12/17/00 16:00 Ole Hansen Shift Summary
  • 53721 12/17/00 08:01 M. Epstein shift summary
  • 53702 12/17/00 01:06 Neyret Compton Beam polarization
  • 53697 12/17/00 00:01 Egle Shift summary 12/16 4-12 pm
  • Saturday

  • 53695 12/16/00 23:42 egle N-Cryolevel low
  • 53675 12/16/00 16:02 T. Chang day shift summary
  • 53669 12/16/00 15:14 T. Chang half-wave plate position changed
  • 53666 12/16/00 14:45 M Jones Keep T1,T2 and T7 triggers
  • 53656 12/16/00 11:09 T. chang event_type distribution changes
  • 53648 12/16/00 07:56 X. Zheng owl shift summary
  • 53628 12/16/00 01:36 Neyret Compton Beam polarization
  • 53622 12/16/00 00:00 Ole Hansen Shift Summary
  • 53620 12/16/00 00:00 R. Michaels R-arm HV on
  • Friday

  • 53600 12/15/00 16:03 gilman calo replay
  • 53598 12/15/00 16:01 T. Chang day shift summary
  • 53597 12/15/00 16:01 T. Chang Q2 nitrogen level is at 10%
  • 53588 12/15/00 12:25 R. Michaels turn on R-arm magnets
  • 53580 12/15/00 09:58 gilman replays of run 1898
  • 53579 12/15/00 09:35 gilman starw efficiency looks good
  • 53576 12/15/00 08:38 gilman fixed things?
  • 53573 12/15/00 07:59 X. Zheng shift summary
  • 53569 12/15/00 06:56 X. Zheng TS1 problem
  • 53555 12/15/00 01:11 Neyret Compton Beam polarization
  • 53548 12/15/00 00:00 R.A.Lindgren summary of runs
  • Thursday

  • 53535 12/14/00 19:36 S.Kozlov LN2 for Q2
  • 53524 12/14/00 16:03 averett day shift summary
  • 53504 12/14/00 07:58 S. Sirca Owl shift summary
  • 53487 12/14/00 01:01 Eugene Moller results
  • Wednesday

  • 53486 12/13/00 23:58 L.Y.Zhu Shift Summary
  • 53483 12/13/00 21:44 L.Y.Zhu beam back
  • 53478 12/13/00 11:51 R. Michaels Trigger work done, R-arm HV off
  • 53472 12/13/00 11:11 gilman Q1 field going to 0
  • 53468 12/13/00 09:41 R. Michaels R-arm HV on
  • 53467 12/13/00 09:01 Wijesooriya Day shift summary 13/12/2000
  • 53462 12/13/00 07:57 S. Sirca Owl shift summary, part II
  • 53453 12/13/00 06:08 S. Sirca Owl shift summary, part I
  • 53449 12/13/00 04:25 S. Sirca Further problems in the injector
  • 53446 12/13/00 03:59 S. Sirca Problem with cavity in injector
  • 53429 12/13/00 00:03 L.Y.Zhu shift summary
  • Tuesday

  • 53418 12/12/00 20:45 L.Y.Zhu FPP HV 7.1.16 trip
  • 53410 12/12/00 17:03 L.Y.Zhu beam off
  • 53405 12/12/00 16:02 wijesooriya day shift summary
  • 53398 12/12/00 12:41 gilman helicity fraction vs run number plot
  • 53386 12/12/00 08:02 S. Sirca Owl shift summary
  • 53375 12/12/00 04:01 S. Sirca Beam dump diffuser problem fixed
  • 53372 12/12/00 01:12 S. Sirca Diffuser flow alarm
  • Monday

  • 53367 12/11/00 23:59 Raue Shift Summary
  • 53366 12/11/00 23:54 Raue Lost beam time
  • 53354 12/11/00 16:04 LeRose Shift summary
  • 53351 12/11/00 15:25 Olivier Dipole field effect on calo
  • 53348 12/11/00 14:27 LeRose Magnet Tests Over
  • 53347 12/11/00 14:23 LeRose Magnet tests
  • 53344 12/11/00 13:36 S. Dieterich emiss shift in calo
  • 53339 12/11/00 12:06 LeRose RF Problems
  • 53338 12/11/00 11:27 LeRose Magnet Tests
  • 53335 12/11/00 10:55 LeRose Magnet Tests
  • 53331 12/11/00 09:47 R. Michaels turn off R-arm HV
  • 53332 12/11/00 09:40 LeRose Right Arm Magnet Tests
  • 53326 12/11/00 08:00 S. Sirca Owl shift summary
  • 53324 12/11/00 07:48 Iodice TDC vs Runnr.
  • 53319 12/11/00 06:22 S. Sirca BPMs A/B x/y do not update
  • 53315 12/11/00 05:36 S. Sirca Master-slave inconsistency
  • 53310 12/11/00 03:44 iodice inconsitencies on paw plots
  • 53305 12/11/00 01:38 S. Sirca What HV 7.0.2 has to do with 7.0.0?
  • 53300 12/11/00 00:05 Raue Shift Summary
  • Sunday

  • 53271 12/10/00 16:01 G.J. Lolos15 shift summary
  • 53248 12/10/00 07:57 S. Sirca Owl shift summary
  • 53243 12/10/00 06:06 kramerk Cryotarget Heater Problems
  • 53227 12/10/00 00:01 Raue summary
  • Saturday

  • 53209 12/09/00 16:07 G.J. Lolos shift summary
  • 53197 12/09/00 10:47 Egle Mauro zeros in nruhits-nrvhits
  • 53194 12/09/00 09:49 Egle Mauro Blocks 136 and 144 all right
  • 53191 12/09/00 08:55 egle, mauro peak at 0 on the react_z spectrum
  • 53187 12/09/00 07:58 S. Sirca Owl shift summary
  • 53186 12/09/00 07:53 S. Sirca FPP HV 7.1.16 trip
  • 53182 12/09/00 07:37 S. Sirca TS1 not responding
  • 53181 12/09/00 07:21 S. Sirca loop_check_fpp.tcl script misbehaves
  • 53176 12/09/00 04:55 S. Sirca FPP HV 7.1.16 trip
  • 53169 12/09/00 04:08 S. Sirca Repeated M3/M4 alarm
  • 53168 12/09/00 03:32 S. Sirca TS1 (halladaq
  • 53166 12/09/00 02:39 S. Sirca M3/M4 alarm cleared
  • 53163 12/09/00 01:02 S. Sirca LHRS D1 M3/M4 over-temperature
  • Friday

  • 53158 12/08/00 23:59 S. Strauch Swing shift summary
  • 53151 12/08/00 21:05 S. Strauch FPP HV 7.1.16 trip
  • 53144 12/08/00 19:07 Jack Left arm HVAC for hut.
  • 53143 12/08/00 16:47 M Jones Down for a/c repairs
  • 53141 12/08/00 14:10 L. Todor Day Shift log
  • 53131 12/08/00 11:04 R. Michaels keep sieve IN on R-arm (DANGER)
  • 53130 12/08/00 11:00 R. Michaels R-arm scint HV turned on
  • 53127 12/08/00 10:02 M JOnes fpp 7.1.16 tripped
  • 53122 12/08/00 08:15 T.Chang T. Chang
  • 53119 12/08/00 08:07 X.Zheng calo calibration
  • 53116 12/08/00 07:52 T. Chang L-arm Q2 helium low
  • 53113 12/08/00 06:34 T.Chang run 1715 is the first LH2 run at 4.8 point
  • 53112 12/08/00 06:15 T. Chang HAC parameters for q^2=4.8
  • 53109 12/08/00 06:05 T. Chang Need to turn on dipole regulator
  • 53108 12/08/00 05:37 T. Chang electron hub temp. raise.
  • 53107 12/08/00 05:29 T.chang HAC parameters after resetting the dipole power
  • 53105 12/08/00 05:25 T.Chang dipole problem
  • 53096 12/08/00 02:32 T. Chang new configuration parameters
  • 53094 12/08/00 00:22 strauch Swing shift summary
  • Thursday

  • 53082 12/07/00 18:08 Jack VDC current readback.
  • 53081 12/07/00 17:18 Strauch Magent screen for sieve setup @ (5.6 GeV/c)^2
  • 53080 12/07/00 16:16 L.Todor Day Shift Activity
  • 53077 12/07/00 13:58 M Jones half wave plate retracted
  • 53068 12/07/00 08:33 gjlolos Run 1699 CODA rates
  • 53064 12/07/00 08:03 T. Chang owl shift summary
  • 53061 12/07/00 07:32 T. Chang electron hub temperature
  • 53060 12/07/00 07:28 T. Chang calo channel 46 almost gone
  • 53057 12/07/00 06:41 T. Chang Argon pressure getting low
  • 53044 12/07/00 03:06 T.Chang Temperature of electron hub
  • 53041 12/07/00 02:32 T.Chang Missing wires in fpp-ru5
  • 53034 12/07/00 00:00 L Bimbot Shift summary
  • Wednesday

  • 53027 12/06/00 21:33 L. Bimbot Increasing temperarture in electron sheld house
  • 53012 12/06/00 16:00 aniol / stra day shift summary
  • 53008 12/06/00 14:36 Jack Resets' display.
  • 53005 12/06/00 14:17 Escoffier Compton results
  • 52989 12/06/00 08:04 S.Kozlov owl shift summary
  • 52985 12/06/00 06:27 S.Kozlov FPP 7.1.16 HV
  • 52984 12/06/00 06:23 S.Kozlov FPP false alarm
  • Tuesday

  • 52964 12/05/00 23:58 D. Margaziot Shift Summary
  • 52954 12/05/00 19:50 Olivier FPP chambers 3 and 4 gas flow read out
  • 52946 12/05/00 16:07 G. M. Urciuo Hall A day shift report
  • 52944 12/05/00 12:01 Olivier calo calibration
  • 52943 12/05/00 11:48 R.Michaels T1 fixed
  • 52941 12/05/00 11:14 x. jiang Change of FPP chamber HV,
  • 52933 12/05/00 09:26 x. jiang magnet screen at 9:30 am
  • 52928 12/05/00 08:03 T. Chang shift summary
  • 52927 12/05/00 08:02 T. Chang gas pressure low
  • 52925 12/05/00 07:17 T.Chang ROC14 loss connection
  • 52911 12/05/00 01:09 T. Chang calo channel 46 bad!
  • 52908 12/05/00 00:01 Aniol shift summary
  • Monday

  • 52898 12/04/00 19:33 Escoffier Compton
  • 52896 12/04/00 19:05 aniol fpp trip
  • 52894 12/04/00 18:54 aniol fpp trip 18:45
  • 52893 12/04/00 18:31 aniol fpp trip at 17:19
  • 52885 12/04/00 16:03 Calarco Shift Summary
  • 52883 12/04/00 15:33 Calarco ROC3
  • 52879 12/04/00 14:03 Jack Argon regulator.
  • 52876 12/04/00 11:18 Calarco FPP HV
  • 52874 12/04/00 11:08 Calarco FPP HV trip
  • 52871 12/04/00 10:55 Calarco FPP HV trip
  • 52869 12/04/00 10:19 Calarco FPP HV trip
  • 52866 12/04/00 09:46 Calarco Target fan speed
  • 52860 12/04/00 08:00 T. Chang shift summary
  • 52853 12/04/00 05:45 T.Chang Argon pressure drop
  • 52846 12/04/00 03:40 T.Chang change to loop_check_fpp. Potential scaler problem.
  • 52841 12/04/00 02:13 T.Chang 8.1.22 FPP trips
  • 52836 12/04/00 01:10 Markowitz 75 microamps
  • 52833 12/04/00 00:44 T. Chang loop_check_fpp.tcl false alarm?
  • Sunday

  • 52830 12/03/00 23:55 Ransome Shift Summary
  • 52814 12/03/00 17:47 Jack Argon system leak.
  • 52809 12/03/00 16:00 Calarco Shift summary
  • 52799 12/03/00 13:50 Escoffier Compton
  • 52795 12/03/00 12:28 Calarco Polarization in other halls
  • 52790 12/03/00 10:35 markowitz half wave plate in
  • 52788 12/03/00 10:24 Calarco Half wave plate
  • 52780 12/03/00 08:01 R. Michaels owl shift summary
  • 52778 12/03/00 06:59 R. Michaels epics screen
  • 52777 12/03/00 06:49 R.Michaels IOCSC10 reboot
  • 52776 12/03/00 06:47 R. Michaels pk/e99007_det kumac prob/soln
  • 52775 12/03/00 06:41 R. Michaels IOCSC10 suspected problems
  • 52770 12/03/00 04:51 R. Michaels HV 8.1.22 trip
  • 52765 12/03/00 03:39 R. Michaels HV trip, end 1583. Then 1584 junk.
  • 52763 12/03/00 03:08 R. Michaels Attempt to study T1,T2,T7. No T1.
  • 52759 12/03/00 01:18 R. Michaels HV 8.1.22 trip
  • 52757 12/03/00 01:05 R. Michaels rocs stop communicating
  • Saturday

  • 52753 12/02/00 23:59 ransome shift summary
  • 52745 12/02/00 20:54 E. Brash Gas Alarm
  • 52739 12/02/00 20:11 Olivier ADC 149 missing
  • 52737 12/02/00 20:09 Olivier Hall C moller
  • 52736 12/02/00 19:51 Olivier Block 35 is retrieved
  • 52730 12/02/00 16:03 Calarco Shift summary
  • 52727 12/02/00 15:31 Escoffier Compton
  • 52718 12/02/00 12:16 Calarco HV trip
  • 52714 12/02/00 11:33 Calarco HV trip
  • 52704 12/02/00 08:00 R. Michaels owl shift summary
  • 52700 12/02/00 07:28 R. Michaels adaqs2 became sluggish
  • 52699 12/02/00 06:53 R. Michaels problems with pk/e99007_det
  • 52696 12/02/00 06:23 R. Michaels calo block 35 missing
  • 52695 12/02/00 05:45 R. Michaels HV 8.1.22 trip, stop run 1557
  • 52691 12/02/00 03:04 R. Michaels BCM fault requires access
  • 52690 12/02/00 02:48 R. Michaels BCM fault
  • 52686 12/02/00 02:11 R. Michaels RF problems
  • 52682 12/02/00 00:49 R. Michaels HV 8.1.22 trip
  • 52678 12/02/00 00:02 S. Strauch Shift summary (swing-shift)
  • Friday

  • 52664 12/01/00 20:03 Strauch Beam profile
  • 52652 12/01/00 18:25 Eugene Moller results
  • 52651 12/01/00 16:27 Saha day shift summary
  • 52649 12/01/00 13:57 mitchell ep energy measurement--preliminary value
  • 52648 12/01/00 13:08 Saha Arc Integral measurement
  • 52647 12/01/00 10:03 mitchell ep beam problems
  • 52645 12/01/00 09:11 Olivier Target low He flow fixed
  • 52642 12/01/00 07:59 R. Michaels owl shift summary
  • 52638 12/01/00 06:23 R. Michaels DAQ hiccup, new calo gate
  • 52636 12/01/00 05:49 R. Michaels Mistake - Calo ADC gate
  • 52635 12/01/00 04:40 R. Michaels Calo ADC Gate -24 nsec
  • 52620 12/01/00 02:19 R. Michaels Beam spot good enough
  • 52609 12/01/00 01:37 R. Michaels spot narrower, but still fat
  • 52603 12/01/00 01:08 R. Michaels turn off R-arm HV
  • 52602 12/01/00 01:04 R. Michaels beam spot, HARP not working